The sum of the first n terms of an arithmetic squence is given by the formula; Sn= n/2 (3n-13)=
i) Find the sum of the first two terms
ii) Find the sum of the first three terms
iii) Find the first three terms of the sequence
iv) How many terms of this sequence are required to give a sum of 85?

Answer by edjones(8007)

iv)
n/2(3n-13)=85
n(3n-13)=170
3n^2-13n-170=0
(n-10)(3n+17)=0
n=10
.
Ed